A quality control inspector found flaws in 13 out of 150 sweaters.
HERE IS THE QUESTION: Suppose the company produces 500 sweaters a day. How many will not have flaws?

Thanks you so much whoever solves this!!!!

Answer by stanbon(75887)

A quality control inspector found flaws in 13 out of 150 sweaters.
HERE IS THE QUESTION: Suppose the company produces 500 sweaters a day. How many will not have flaws
---------
Proportion:
x/500 = 137/150
x = (137/150)*500
x = 456.667
--------------------
Rounding down you get:
# without flaws is 456 out of 500
===================
Cheers,
Stan H.